# TartanGround: A Large-Scale Dataset for Ground Robot Perception and Navigation

### v1

@@ Abstract
We present TartanGround, a large-scale, multi-modal dataset to advance the perception and autonomy of ground robots operating in diverse environments. This dataset, collected in various photorealistic simulation environments includes multiple RGB stereo cameras for 360-degree coverage, along with depth, optical flow, stereo disparity, LiDAR point clouds, ground truth poses, semantic segmented images, and occupancy maps with semantic labels. Data is collected using an integrated automatic pipeline, which generates trajectories mimicking the motion patterns of various ground robot platforms, including wheeled and legged robots. We collect 910 trajectories across 70 environments, resulting in 1.5 million samples. Evaluations on occupancy prediction and SLAM tasks reveal that state-of-the-art methods trained on existing datasets struggle to generalize across diverse scenes. TartanGround can serve as a testbed for training and evaluation of a broad range of learning-based tasks, including occupancy prediction, SLAM, neural scene representation, perception-based navigation, and more, enabling advancements in robotic perception and autonomy towards achieving robust models generalizable to more diverse scenarios. The dataset and codebase for data collection will be made publicly available upon acceptance.
Webpage: https://tartanair.org/tartanground
@@ I Introduction
To address the need for comprehensive data resources, we introduce TartanGround, a large-scale simulation dataset designed to support perception and navigation tasks for ground robots across diverse environments. TartanGround includes data from over 70 diverse, challenging photorealistic environments, offering multi-modal sensor data such as multiple stereo RGB-D images covering a 360°, semantic labels, LiDAR point clouds, semantic occupancy maps, and ground truth poses. We show that the existing occupancy prediction methods, solely trained on data from the autonomous driving domain, do not generalize to environments of mobile robot operations such as forests. Moreover, we evaluate various algorithms and find that they struggle in challenging scenarios of low visibility and heavy occlusions. To summarize, the main contributions of this paper are: (1) A large dataset with over 1.5 million samples collected across diverse environments with precise ground truth labels mimicking different ground robot motion patterns, including wheeled (omnidirectional and differential-drive), and legged robots (quadrupedal), (2) an automatic data collection pipeline and (3) evaluation of two tasks, occupancy prediction and highlighting the limitations of models trained on existing datasets.
@@ II Related Work
| Dataset | Env | Motion Pattern | Stereo | Depth | LiDAR | Semantics | Occupancy | Pose | Multi-cam | Seq Num | Samples |
| Semantic KITTI [21] | Urban | Car | ✔ | ✘ | ✔ | ✔ | ✔ | ✔ | ✘ | 22 | 43 K |
| Occ3D-Waymo [2, 22] | Urban | Car | ✘ | ✘ | ✔ | ✔ | ✔ | ✔ | 5 | 1000 | 200 K |
| Occ3D-nuScenes [4, 22] | Urban | Car | ✘ | ✘ | ✔ | ✔ | ✔ | ✔ | 6 | 1000 | 40 K |
| RELLIS-3D [5] | Off-road | Car | ✔ | ✔ | ✔ | ✔ | ✘ | ✔ | ✘ | 5 | 6.2 K |
| RUGD [6] | Off-road | Car | ✘ | ✘ | ✘ | ✔ | ✘ | ✘ | ✘ | 18 | 7.5 K |
| TartanDrive [8] | Off-road | Car | ✔ | ✔ | ✘ | ✘ | ✘ | ✔ | ✘ | 630 | 184 K |
| WildScenes [10] | Off-road | Hand | ✘ | ✘ | ✔ | ✔ | ✘ | ✔ | ✘ | 5 | 9.6 K |
| WildOcc [25] | Off-road | Car | ✔ | ✔ | ✔ | ✔ | ✔ | ✔ | ✘ | 5 | 10 K |
| GOOSE [7] | Off-road | Car | ✘ | ✘ | ✔ | ✔ | ✔ | ✔ | 4 | 356 | 10 K |
| Scannet [11] | Indoor | Hand | ✘ | ✔ | ✘ | ✔ | ✔ | ✔ | ✘ | 1500 | 2.5 M |
| SUN RGBD [23] | Indoor | Hand | ✘ | ✔ | ✘ | ✔ | ✘ | ✘ | ✘ | - | 10 K |
| GrandTour [27] | Mix | Legged, Wheeled | ✔ | ✔ | ✔ | ✘ | ✘ | ✔ | 3 | 100 | 100 K |
| TartanAir [14] | Mix | Random | ✔ | ✔ | ✘ | ✔ | ✘ | ✔ | ✘ | 1037 | 400 k |
| TartanAir-V2 [26] | Mix | Random | ✔ | ✔ | ✔ | ✔ | ✔ | ✔ | 6 | 1110 | 1.4 M |
| TartanGround | Mix | Legged, Wheeled | ✔ | ✔ | ✔ | ✔ | ✔ | ✔ | 6 | 910 | 1.5 M |
@@ III-A Features
TartanGround consists of 70 realistic simulation environments from TartanAir-V2 [26], covering diverse scenarios from structured urban to large-scale unstructured outdoor environments (Fig. 2). The environments have been developed with Unreal Engine, and the data is collected using the AirSim [28] plugin. This setup allows for rendering photorealistic scenes with high fidelity and makes it possible to have dynamic lighting, adverse weather effects, dynamic objects, and seasonal changes. In each of the environments, we collect data from multiple sampled trajectories with diverse motion patterns mimicking real-world ground robots. To achieve full 360 coverage, we use 6 stereo RGB cameras (front, left, right, back, top, bottom), each with a of 90. Each camera is synchronized with accurate ground truth poses and also records depth and semantic segmentation images. During post-processing, we generate additional ground truth data, including IMU data, optical flow, stereo disparity, LiDAR point clouds, and semantic occupancy maps. In total, we collect 910 trajectories (441 omni-wheeled, 226 diff-wheeled, and 243 legged), each with 600 to 8000 samples, resulting in a total dataset size of approximately 15 TB, 1.5 million samples, and 18 million RGB images. This large-scale, multi-modal dataset is designed to support a wide range of robotic perception and navigation tasks, benefiting the community in establishing new benchmarks for generalizable learning-based methods.
@@ IV-B1 Testing Environments and Baselines
In this section, we show a small-scale evaluation of three state-of-the-art VO/SLAM algorithms on four challenging trajectories from four different environments. The trajectory from the Forest environment contains occlusion from tall grass and bushes, which are commonly presented in real-world off-road scenes. CastleFortress is a large scene with indoor-outdoor transitions, bringing dramatic illumination change. WaterMillDay has a rushing creek with water splash. ModUrbanCity contains a narrow dark stair that lacks good visual features. Our baseline consists of a classic geometry-based SLAM algorithm ORB-SLAM3 [33], a learning-based monocular odometry DPVO [34], and a stereo VO model MACVO [35]. We use relative translation error (, ) and relative rotation error (, ) to evaluate the results.
@@ IV-D Neural Scene Representation
The photorealism of our dataset makes it well-suited for advancing research in neural scene representation techniques, such as Gaussian splatting [41] and NeRFs [42], as well as neural SLAM methods [43, 44, 45]. Its large-scale scenarios, with dynamic lighting and adverse weather conditions, offer challenging test cases for novel view synthesis and robust scene reconstruction. As shown in Fig. 7, novel view synthesis using Gaussian splatting trained on trajectories from the Rome and CoalMine environments demonstrates the potential of TartanGround for high-fidelity neural rendering.
@@ Acknowledgements
The authors would like to thank Jannick Schröer for generating the Gaussian splatting renderings.

### v2

@@ Abstract
We present TartanGround, a large-scale, multi-modal dataset to advance the perception and autonomy of ground robots operating in diverse environments. This dataset, collected in various photorealistic simulation environments includes multiple RGB stereo cameras for 360-degree coverage, along with depth, optical flow, stereo disparity, LiDAR point clouds, ground truth poses, semantic segmented images, and occupancy maps with semantic labels. Data is collected using an integrated automatic pipeline, which generates trajectories mimicking the motion patterns of various ground robot platforms, including wheeled and legged robots. We collect 878 trajectories across 63 environments, resulting in 1.44 million samples. Evaluations on occupancy prediction and SLAM tasks reveal that state-of-the-art methods trained on existing datasets struggle to generalize across diverse scenes. TartanGround can serve as a testbed for training and evaluation of a broad range of learning-based tasks, including occupancy prediction, SLAM, neural scene representation, perception-based navigation, and more, enabling advancements in robotic perception and autonomy towards achieving robust models generalizable to more diverse scenarios. The dataset and codebase are available on the webpage: https://tartanair.org/tartanground
@@ I Introduction
To address the need for comprehensive data resources, we introduce TartanGround, a large-scale simulation dataset designed to support perception and navigation tasks for ground robots across diverse environments. TartanGround includes data from over 63 diverse, challenging photorealistic environments, offering multi-modal sensor data such as multiple stereo RGB-D images covering a 360°, semantic labels, LiDAR point clouds, semantic occupancy maps, and ground truth poses. We show that the existing occupancy prediction methods, solely trained on data from the autonomous driving domain, do not generalize to environments of mobile robot operations such as forests. Moreover, we evaluate various algorithms and find that they struggle in challenging scenarios of low visibility and heavy occlusions. To summarize, the main contributions of this paper are: (1) A large dataset with over 1.44 million samples collected across diverse environments with precise ground truth labels mimicking different ground robot motion patterns, including wheeled (omnidirectional and differential-drive), and legged robots (quadrupedal), (2) an automatic data collection pipeline and (3) evaluation of two tasks, occupancy prediction and highlighting the limitations of models trained on existing datasets.
@@ II Related Work
| Dataset | Type | Env | Motion Pattern | Stereo | Depth | LiDAR | Semantics | Occupancy | Pose | Multi-cam | Seq Num | Samples |
| Semantic KITTI [21] | Real | Urban | Car | ✔ | ✘ | ✔ | ✔ | ✔ | ✔ | ✘ | 22 | 43 K |
| Occ3D-Waymo [2, 22] | Real | Urban | Car | ✘ | ✘ | ✔ | ✔ | ✔ | ✔ | 5 | 1000 | 200 K |
| Occ3D-nuScenes [4, 22] | Real | Urban | Car | ✘ | ✘ | ✔ | ✔ | ✔ | ✔ | 6 | 1000 | 40 K |
| RELLIS-3D [5] | Real | Off-road | Car | ✔ | ✔ | ✔ | ✔ | ✘ | ✔ | ✘ | 5 | 6.2 K |
| RUGD [6] | Real | Off-road | Car | ✘ | ✘ | ✘ | ✔ | ✘ | ✘ | ✘ | 18 | 7.5 K |
| TartanDrive [8] | Real | Off-road | Car | ✔ | ✔ | ✘ | ✘ | ✘ | ✔ | ✘ | 630 | 184 K |
| WildScenes [10] | Real | Off-road | Hand | ✘ | ✘ | ✔ | ✔ | ✘ | ✔ | ✘ | 5 | 9.6 K |
| WildOcc [25] | Real | Off-road | Car | ✔ | ✔ | ✔ | ✔ | ✔ | ✔ | ✘ | 5 | 10 K |
| SynPhoRest [27] | Syn. | Off-road | Random | ✘ | ✔ | ✔ | ✔ | ✘ | ✘ | ✘ | X | 3.1 K |
| GOOSE [7] | Real | Off-road | Car | ✘ | ✘ | ✔ | ✔ | ✔ | ✔ | 4 | 356 | 10 K |
| GOOSE-Ex [28] | Real | Off-road | Tracked, Legged | ✘ | ✘ | ✔ | ✔ | ✔ | ✔ | 4 | 100 | 5 K |
| Scannet [11] | Real | Indoor | Hand | ✘ | ✔ | ✘ | ✔ | ✔ | ✔ | ✘ | 1500 | 2.5 M |
| GrandTour [29] | Real | Mix | Legged, Wheeled | ✔ | ✔ | ✔ | ✘ | ✘ | ✔ | 3 | 100 | 100 K |
| TartanAir [14] | Syn. | Mix | Random | ✔ | ✔ | ✘ | ✔ | ✘ | ✔ | ✘ | 1037 | 400 k |
| TartanAir-V2 [26] | Syn. | Mix | Random | ✔ | ✔ | ✔ | ✔ | ✔ | ✔ | 6 | 1110 | 1.4 M |
| TartanGround | Syn. | Mix | Legged, Wheeled | ✔ | ✔ | ✔ | ✔ | ✔ | ✔ | 6 | 878 | 1.4 M |
@@ III-A Features
TartanGround consists of 63 realistic simulation environments from TartanAir-V2 [26], covering diverse scenarios from structured urban to large-scale unstructured outdoor environments (Fig. 2). The environments have been developed with Unreal Engine 4, and the data is collected using the AirSim [30] plugin. This setup allows for rendering photorealistic scenes with high fidelity and makes it possible to have dynamic lighting, adverse weather effects, dynamic objects, and seasonal changes. In each of the environments, we collect data from multiple sampled trajectories with diverse motion patterns mimicking real-world ground robots. To achieve full 360 coverage, we use 6 stereo RGB cameras (front, left, right, back, top, bottom), each with a of 90. Each camera is synchronized with accurate ground truth poses and also records depth and semantic segmentation images. During post-processing, we generate additional ground truth data, including IMU data, optical flow, stereo disparity, LiDAR point clouds, and semantic occupancy maps. In total, we collect 878 trajectories (440 omni-wheeled, 198 diff-wheeled, and 240 legged), each with 600 to 8000 samples, resulting in a total dataset size of approximately 15 TB, 1.44 million samples, and 17.3 million RGB images. This large-scale, multi-modal dataset is designed to support a wide range of robotic perception and navigation tasks, benefiting the community in establishing new benchmarks for generalizable learning-based methods.
@@ IV-B1 Testing Environments and Baselines
In this section, we show a small-scale evaluation of three state-of-the-art VO/SLAM algorithms on four challenging trajectories from four different environments. We use the front camera for these experiments. The trajectory from the Forest environment contains occlusion from tall grass and bushes, which are commonly presented in real-world off-road scenes. CastleFortress is a large scene with indoor-outdoor transitions, bringing dramatic illumination change. WaterMillDay has a rushing creek with water splash. ModUrbanCity contains a narrow dark stair that lacks good visual features. Our baseline consists of a classic geometry-based SLAM algorithm ORB-SLAM3 [35], a learning-based monocular odometry DPVO [36], and a stereo VO model MACVO [37]. We use relative translation error (, ) and relative rotation error (, ) to evaluate the results.
@@ IV-D Neural Scene Representation
The photorealism of our dataset makes it well-suited for advancing research in neural scene representation techniques, such as Gaussian splatting [43] and NeRFs [44], as well as neural SLAM methods [45, 46, 47, 48]. Its large-scale scenarios, with dynamic lighting and adverse weather conditions, offer challenging test cases for novel view synthesis and robust scene reconstruction. As shown in Fig. 7, novel view synthesis using Gaussian splatting trained on trajectories from the Rome and CoalMine environments demonstrates the potential of TartanGround for high-fidelity neural rendering.
@@ Acknowledgements
This work was supported by the Luxembourg National Research Fund (Ref. 18990533) and the Swiss National Science Foundation (Ref. 200021E_229503). This work used Bridges-2 at PSC through allocation cis220039p from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program which is supported by NSF grants 2138259, 2138286, 2138307, 2137603, and 2138296. The authors would like to thank Jannick Schröer for generating the Gaussian splatting renderings.